A salt solution containing ferrous ion was titrated with (9.80000x10^-3) M KMnO4 solution. The mean of 3 acceptable, corrected titration values was (1.400x10^1) mL. Calculate the moles of ferrous ion reacting in each titration portion.
1
Expert's answer
2020-01-13T03:32:18-0500

5Fe2++MnO4-+8H+"\to" 5Fe3++Mn2++4H2O

1.400x10^1(mL)=0.014L

n(KMnO4)=9.80000x10^-3(M)*0.014(L)=1.372*10^-4(moles)

n(Fe2+)=1.372*10^-4(moles)*5=6.86*10^-4(moles).